OpenStudy (anonymous):

What is the approximate area of the shaded region? A. 178.5 square centimeters B. 100 square centimeters C. 78.5 square centimeters D. 21.5 square centimeters

OpenStudy (anonymous):

Picture?

OpenStudy (anonymous):

Area of Circle: pi*r^2 Area of Square: side^2 Square: 10*10=100 Circle: 3.14*25=78.5 100-78.5=21.5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no prob.